at a movie theater the adult ticket price was 8$ &a the child ticket price was 6$. for a certain movie, 200 tickets were sold & 1440$ was collected how may adult tickets were sold
Answer by Boreal(15235)

A+C=200, the ticket number equation
8A+6C=1440, the cost equation
multiply the first by -6 to eliminate C
-6A-6C=-1200
2A=240
A=120 adults@$8=$960
C=80 child @$6=$480
$1440.